OpenStudy (anonymous):

if f(x)=8x and g(x)=9x-6 find (fg)(x)

OpenStudy (zzr0ck3r):

\[f(x) = 8x\\g(x) = 9x-6\\so\\(fg)(x)=f(x)*g(x) = (8x)(9x-6)=(8x)(9x)-(8x)(6)=72x^2-48x\]